Database last updated on Wed, 04 Feb 2026 07:40:05 +0000 by localhost
Instrument constant changes for the week before 2023312 (1 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 14 rows)
Measurements and tests summary.
| _jday | _loc | _serial |
| 312 | TORONTO | #187 |
(0 sec)
Dead time (1 sec)
Nothing to report.
Standard lamp R6 (1 sec)
Nothing to report.
Standard lamp R5 (1 sec)
Nothing to report.
Number of days since last dome/window cleaning (0 sec, 3 rows)
| serial | days |
| 187 | more than a year |
| 8 | more than a year |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 187 | 18 | 4 | 7 | 6 | 3 | 6 | 4 | 3 | 3 |
Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)
No ozone measurements to display.
Unsorted comments (1 sec, 10159 rows)
The END (0 sec, 2 rows)